Suppose a viscous oil, whose flow is in the laminar regime is to be pumped through a 10 cm
diameter horizontal pipe over a distance of 15km at a rate of s/ 10 m
−3 3
. Viscosity of the oil is
03.0 poise. What is the required pressure drop to maintain such a flow?
1
Expert's answer
2020-03-26T14:32:52-0400

Suppose a viscous oil, whose flow is in the laminar regime is to be pumped through a 10 cm diameter horizontal pipe over a distance of 15km at a rate of "10^{-3}\ m^3/s" . Viscosity of the oil is 3.0 poise. What is the required pressure drop to maintain such a flow? 

Solution

For a viscous oil at 20°C take "\rho=100\ kg/m^3 ." Given that "\mu=3 \ poise=0.3\ \dfrac{N\cdot s}{m^2},"

"D=10 \ cm=0.1\ m, L=15\ km=15000\ m, Q=10^{-3}\ m^3/s."

Calculate the pipe velocity and Reynolds number:


"V={Q \over A}={Q \over (\pi D^2)/4}={10^{-3}\ m^3/s \over (\pi (0.1\ m)^2)/4}\approx0.1273\ m/s"

"Re={\rho VD \over \mu}={1000\dfrac{kg}{m^3}\cdot0.1273\dfrac{m}{s}\cdot0.1\ m\over 0.3\dfrac{N\cdot s}{m^2}}=42.4<2100"

The flow is laminar flow. Find the required pressure drop to maintain such a flow


"\Delta p=p_2-p_1={128\mu LQ \over \pi D^4}"

"\Delta p={128\cdot0.3\ \dfrac{N\cdot s}{m^2}\cdot15000\ m\cdot10^{-3}\ \dfrac{m^3}{s} \over \pi (0.1\ m)^4 }="

"=183346.5 \text{Pa}=1.833465\times10^5 \text{Pa}=0.183\ \text{MPA}"